OpenStudy (anonymous):

Find the quotient. 42j^4k^2 ÷ (-3j^3k) -14jk 14jk -14j²k

OpenStudy (anonymous):

anyone know

OpenStudy (ash2326):

\[\large \frac{42j^4k^2}{-3j^3k}\] Let's cancel the common terms \[\large \frac{42j^{\cancel{4}1}k^{\cancel{2}1}}{-3\cancel{j^3}\cancel k}\] \[\large \frac{42 jk}{-3}\] \[-\large \frac{\cancel{42}14 jk}{\cancel{3}}\] so we get finally \[\huge -14jk\]

OpenStudy (anonymous):

ok thank you

OpenStudy (ash2326):

Do you understand?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

ya i do